variable cost per unit produced and sold

Kubin Company’s relevant range of production is 18,000 to 22,000 units. When it produces and sells 20,000 units, its average costs per unit are as follows:

 


Amount per Unit
Direct materials$7.00
Direct labor$4.00
Variable manufacturing overhead$1.50
Fixed manufacturing overhead$5.00
Fixed selling expense$3.50
Fixed administrative expense$2.50
Sales commissions$1.00
Variable administrative expense$0.50

 

Required:

1. If 18,000 units are produced and sold, what is the variable cost per unit produced and sold?

2. If 22,000 units are produced and sold, what is the variable cost per unit produced and sold?

3. If 18,000 units are produced and sold, what is the total amount of variable cost related to the units produced and sold?

4. If 22,000 units are produced and sold, what is the total amount of variable cost related to the units produced and sold?

5. If 18,000 units are produced, what is the average fixed manufacturing cost per unit produced?

6. If 22,000 units are produced, what is the average fixed manufacturing cost per unit produced?

7. If 18,000 units are produced, what is the total amount of fixed manufacturing overhead incurred to support this level of production?

8. If 22,000 units are produced, what is the total amount of fixed manufacturing overhead incurred to support this level of production?
Answer #1
Concepts and reason

Product Cost: Product cost can be defined as the total cost incurred during the manufacturing of a product that includes direct labor, direct material and manufacturing overheads.

Absorption Costing: It refers to one of the methods to compute the cost of the manufacturing product. While computing the product cost, this method takes total cost of the product that includes variable overhead manufacturing cost as well as fixed overhead manufacturing cost.

Period Cost: It is the cost that is neither an essential part of the manufacturing process nor directly associated with the manufacturing cost.

Fundamentals

Direct Material: It can be defined as those supplies and materials that are used for manufacturing the product and can be directly associated with the product.

Direct Labor: Direct labor can be defined as the personnel that are engaged directly in the manufacturing of products rather than its maintenance, administration, or other support services.

Variable overhead: It can be defined as a manufacturing cost that varies as per the level of production. This is used to determine future costs and expenditure of a manufacturing company. It is also used to determine the lowest possible price at which the company’s product can be sold. Example: electricity consumed by machinery.

Fixed Overhead: It refers to the cost that is to be incurred, irrespective of the level of production. They do not vary due to the change in the level of an organization’s activities. These costs are required to operate the business. Example: rent of a building, insurance, etc.

Variable Administrative Expenses: These are operating expenses other than the manufacturing costs. Some of the examples of variable administrative expenses are a commission, freight, salaries, etc.

Fixed Administrative Expenses: These refer to the operating expenses that do not change, irrespective of the level of goods manufactured. Some of the examples are rent and insurance.

Sales Commission: It can be defined as an additional income received by an employee if he meets as well as exceeds the minimum level of threshold set by the employer.

Selling expenses: These can be defined as the costs associated with selling and distribution, marketing, advertising, website maintenance, and social media expenses.

(1)

Compute the product cost incurred to make 20,000 units by K Company as shown below:

ProductCost=Directmaterial+DirectLabour+Variablemanufacturingoverhead+FixedManufacturingOverhead=$140,000+$80,000+$30,000+$100,000=$350,000\begin{array}{c}\\{\rm{Product Cost}} = {\rm{Direct material}} + {\rm{Direct Labour}} + {\rm{Variable manufacturing }}\\\\{\rm{ overhead}} + {\rm{FixedManufacturing Overhead }}\\\\ = {\rm{\$ 140,000}} + {\rm{\$ 80,000}} + {\rm{\$ 30,000}} + {\rm{\$ 100,000}}\\\\ = {\rm{\$ 350,000}}\\\end{array}

Therefore, the product cost is $350,000\$ 350,000 .

Working Notes:

Compute the direct material as shown below:

DirectMaterials=NumberofUnits×AverageCostPerunit=20,000units×$7perunit=$140,000\begin{array}{c}\\{\rm{Direct Materials}} = {\rm{Number of Units}} \times {\rm{Average Cost Per unit}}\\\\ = 20,000\,{\rm{units}} \times \$ 7\,{\rm{per}}\,{\rm{unit}}\\\\ = \$ 140,000\\\end{array}

Hence, the direct material is $140,000\$ 140,000 .

Compute the direct labor as shown below:

DirectLabor=NumberofUnits×AverageCostPerunit=20,000units×$4perunit=$80,000\begin{array}{c}\\{\rm{Direct Labor}} = {\rm{Number of Units}} \times {\rm{Average Cost Per unit}}\\\\ = 20,000\,{\rm{units}} \times \$ 4\,{\rm{per}}\,{\rm{unit}}\\\\ = \$ 80,000\\\end{array}

Hence, direct labor is $80,000\$ 80,000 .

Compute the variable manufacturing overhead as shown below:

VariableManufacturingOverhead=NumberofUnits×AverageCostPerunit=20,000units×$1.5perunit=$30,000\begin{array}{c}\\{\rm{Variable Manufacturing Overhead}} = {\rm{Number of Units}} \times {\rm{Average Cost Per unit}}\\\\ = 20,000\,{\rm{units \times }}\$ 1.5\,{\rm{per}}\,{\rm{unit}}\\\\ = \$ 30,000\\\end{array}

Hence, variable manufacturing overhead is $30,000\$ 30,000 .

Compute the fixed manufacturing overhead as shown below:

FixedManufacturingOverhead=NumberofUnits×AverageCostPerunit=20,000units×$5perunit=$100,000\begin{array}{c}\\{\rm{Fixed Manufacturing Overhead = Number of Units \times Average Cost Per unit}}\\\\ = 20,000\,{\rm{units }} \times {\rm{ }}\$ 5\,{\rm{per unit}}\\\\ = {\rm{\$ 100,000}}\\\end{array}

Hence, the fixed manufacturing overhead is $100,000\$ 100,000 .

(2)

The period cost incurred to make 20,000 units by K Company is computed as shown below:

Periodcosts=Fixedsellingexpenses+Fixedadministrativeexpenses+Variableadministrativeexpenses+SalesCommission=$70,000+$50,0000+$10,000+$20,000=$150,000\begin{array}{c}\\{\rm{Period costs}} = {\rm{Fixed selling expenses}} + {\rm{Fixed administrative expenses}}\\\\ + {\rm{Variable administrative expenses}} + {\rm{Sales Commission}}\\\\ = \$ 70,000 + {\rm{\$ 50,0000}} + {\rm{\$ 10,000}} + \$ 20,000\\\\ = {\rm{\$ 150,000}}\\\end{array}

Therefore, the period cost incurred to produce 20,000 units is $150,000.

Working Notes:

The fixed selling expenses for 20,000 units is computed as follows:

Fixedsellingexpenses=Units×Costperunit=20,000units×$3.5perunit=$70,000\begin{array}{c}\\{\rm{Fixed selling expenses}} = {\rm{Units}} \times {\rm{Cost per unit}}\\\\ = 20,000\,{\rm{units}} \times \$ 3.5\,{\rm{per}}\,{\rm{unit}}\\\\ = \$ 70,000\\\end{array}

Hence, the fixed selling expenses for 20,000 units is $70,000\$ 70,000 .

The fixed administrative expenses for 20,000 units is computed as follows:

Fixedadministrativeexpenses=Units×Costperunit=20,000units×$2.5perunit=$50,000\begin{array}{c}\\{\rm{Fixed administrative expenses}} = {\rm{Units}} \times {\rm{Cost per unit}}\\\\ = 20,000\,{\rm{units}} \times \$ 2.5\,{\rm{per unit}}\\\\ = \$ 50,000\\\end{array}

Hence, the fixed administrative expenses for 20,000 units is $50,000\$ 50,000 .

The variable administrative expenses for 20,000 units is computed as follows:

Variableadministrativeexpenses=Units×Costperunit=20,000units×$0.5perunit=$10,000\begin{array}{c}\\{\rm{Variable administrative expenses}} = {\rm{Units}} \times {\rm{Cost per unit}}\\\\ = 20,000\,{\rm{units}} \times \$ 0.5\,{\rm{per unit}}\\\\ = \$ 10,000\\\end{array}

Hence, the variable administrative expenses for 20,000 units is $10,000\$ 10,000 .

The sales commission expenses for 20,000 units is computed as follows:

Salescommission=Units×Costperunit=20,000units×$1perunit=$20,000\begin{array}{c}\\{\rm{Sales commission}} = {\rm{Units}} \times {\rm{Cost per unit}}\\\\ = 20,000\,{\rm{units}} \times \$ 1\,{\rm{per unit}}\\\\ = \$ 20,000\\\end{array}

Hence, the sales commission for 20,000 units is $20,000\$ 20,000 .

(3)

Compute the product cost incurred to make 22,000 units by K Company:

ProductCost=Directmaterial+DirectLabor+Variablemanufacturingoverhead+FixedManufacturingOverhead=$154,000+$88,000+$33,000+$100,000=$375,000\begin{array}{c}\\{\rm{Product Cost}} = {\rm{Direct material}} + {\rm{Direct Labor}} + {\rm{Variable manufacturing }}\\\\{\rm{ overhead}} + {\rm{FixedManufacturing Overhead }}\\\\ = {\rm{\$ 154,000}} + {\rm{\$ 88,000}} + {\rm{\$ 33,000}} + {\rm{\$ 100,000}}\\\\ = {\rm{\$ 375,000}}\\\end{array}

Therefore, the product cost incurred to produce 22,000 units is $375,000\$ 375,000 .

Working Notes:

The cost of direct materials is computed as shown below:

DirectMaterials=NumberofUnits×AverageCostPerunit=22,000units×$7perunit=$154,000\begin{array}{c}\\{\rm{Direct Materials}} = {\rm{Number of Units}} \times {\rm{Average Cost Per unit}}\\\\ = 22,000\,{\rm{units}} \times \$ 7\,{\rm{per unit}}\\\\ = {\rm{\$ 154,000}}\\\end{array}

Hence, the direct materials are worth $154,000\$ 154,000 .

The cost of direct labor is computed as shown below:

DirectLabour=NumberofUnits×AverageCostPerunit=22,000units×$4perunit=$88,000\begin{array}{c}\\{\rm{Direct Labour}} = {\rm{Number of Units }} \times {\rm{ Average Cost Per unit}}\\\\ = 22,000\,{\rm{units}} \times \$ 4\,{\rm{per}}\,{\rm{unit}}\\\\ = \$ 88,000\\\end{array}

Hence, direct labor costs $88,000\$ 88,000 .

Compute the variable manufacturing overhead as shown below:

VariableManufacturingOverhead=NumberofUnits×AverageCostPerunit=22,000units×$1.5perunit=$33,000\begin{array}{c}\\{\rm{Variable Manufacturing Overhead}} = {\rm{Number of Units}} \times {\rm{Average Cost Per unit}}\\\\ = 22,000\,{\rm{units}} \times \$ 1.5\,{\rm{per}}\,{\rm{unit}}\\\\ = \$ 33,000\\\end{array}

Hence, the variable manufacturing overhead is $33,000\$ 33,000 .

Compute the fixed manufacturing overhead as shown below:

FixedManufacturingOverhead=NumberofUnits×AverageCostPerunit=20,000units×$5perunit=$100,000\begin{array}{c}\\{\rm{Fixed Manufacturing Overhead}} = {\rm{Number of Units}} \times {\rm{Average Cost Per unit}}\\\\ = 20,000\,{\rm{units}} \times \$ 5\,{\rm{per}}\,{\rm{unit}}\\\\ = \$ 100,000\\\end{array}

Hence, the fixed manufacturing overhead is $100,000\$ 100,000 .

(4)

Compute the period cost incurred to make 18,000 units by K Company:

PeriodCost=FixedSellingExpenses+FixedAdministrativeExpenses+VariableAdministrativeexpenses+SalesCommission=$70,000+$50,000+$9,000+$18,000=$147,000\begin{array}{c}\\{\rm{Period Cost}} = {\rm{Fixed Selling Expenses}} + {\rm{Fixed Administrative Expenses}} + \\\\{\rm{ Variable Administrative expenses + Sales Commission }}\\\\ = {\rm{\$ 70,000 + \$ 50,000 + \$ 9,000 + \$ 18,000}}\\\\ = {\rm{\$ 147,000}}\\\end{array}

Therefore, the period cost is $147,000\$ 147,000 .

Working Notes

The fixed selling expenses are computed as follows:

FixedSellingExpenses=NumberofUnits×AverageCostPerunit=20,000units×$3.5perunit=$70,000\begin{array}{c}\\{\rm{Fixed Selling Expenses = Number of Units }} \times {\rm{ Average Cost Per unit}}\\\\{\rm{ = }}20,000\,{\rm{units}} \times \$ 3.5\,{\rm{per}}\,{\rm{unit}}\\\\ = \$ 70,000\\\end{array}

Hence, the fixed selling expenses are $70,000\$ 70,000 .

The fixed administrative expenses are computed as follows:

FixedAdministrativeExpenses=NumberofUnits×AverageCostPerunit=20,000units×$2.5perunit=$50,000\begin{array}{c}\\{\rm{Fixed Administrative Expenses = Number of Units }} \times {\rm{ Average Cost Per unit}}\\\\{\rm{ = }}20,000\,{\rm{units}} \times \$ 2.5\,{\rm{per}}\,{\rm{unit}}\\\\ = \$ 50,000\\\end{array}

Hence, the fixed administrative expenses are $50,000\$ 50,000 .

The variable administrative expenses are computed as follows:

VariableAdministrativeExpenses=NumberofUnits×AverageCostPerunit=18,000units×$0.5perunit=$9,000\begin{array}{c}\\{\rm{Variable Administrative Expenses = Number of Units \times Average Cost Per unit}}\\\\ = 18,000\,{\rm{units }} \times {\rm{ }}\$ 0.5\,{\rm{per}}\,{\rm{unit}}\\\\ = {\rm{\$ 9,000}}\\\end{array}

Hence, the variable administrative expenses are $9,000\$ 9,000 .

The sales commission is computed as follows:

SalesCommission=NumberofUnits×AverageCostPerunit=18,000units×$1perunit=$18,000\begin{array}{c}\\{\rm{Sales Commission = Number of Units }} \times {\rm{ Average Cost Per unit}}\\\\{\rm{ = }}18,000\,{\rm{units}} \times \$ 1\,{\rm{per}}\,{\rm{unit}}\\\\ = \$ 18,000\\\end{array}

Hence, the sales commission amounts to $18,000\$ 18,000 .

Ans: Part 1

The total product cost for producing 20,000 units is $350,000.

Part 2

The total period cost for producing 20,000 units is $150,000.

Part 3

The total product cost for producing 22,000 units is $375,000.

Part 4

The total period cost for producing 18,000 units is $147,000.
Answer #2

  

Total Fixed Manufacturing overhead = Current levels ofUnits * FMOH per unit

= $5.00* 20,000 = 100,000

Current Level of production in Unit Variable cost Per Unit Direct Material (DM) Direct Labor (DL) Variable Manufacturing Over5 18,000 Average Fixed Cost Per Unit (100000/18000) 6 22000 Average Fixed Cost Per Unit (100000/22000) 7 18,000 Total Amount